The molar ratio of insert DNA to vector DNA used in a ligation is generally about 2:1. Using a vector that is 8 kb and an insert that is 4 kb, how much insert would you use if you used 20 ng of vector?
Answer #1

Required mass insert (ng) = deisred insert/vector molar ratio * mass of vector (ng)*ratio of insert to vector length

Ratio of insert to vector length = 4/8 = 1/2

mass of vector = 20 ng

Insert/Vector molar ratio : 2/1

Requires mass (ng) = 2/1*20*1/2

= 20 ng

Hence we should use 20ng of insert